# rerankers: A Lightweight Python Library to Unify Ranking Methods

### v1

@@ 1 Introduction
Over time, there has been an increasing number of neural re-ranking methods. Most early neural re-ranking models were cross-encoders,  [12], framing the task as a binary sentence classification task and using a pre-trained encoder model such as BERT [4]. However, many new methods have appeared, such as MonoT5 [11], leveraging sequence-to-sequence models such as T5c̃itet5 rather than BERT-like models, or ColBERT-based methods, which repurpose late-interaction retrieval mdoels such as ColBERTv2 [18] as re-ranking models. Other work also explores different paradigms, such as listwise re-ranking using sequence-to-sequence models, where the model is not tasked with outputting a score for a [query, document] pairs, but rather, to directly output a list of documents ordered by their relevance to the given query [21]. Work such as FlashRank [3], has focused on optimised methods to speed up re-ranking when performed on weaker inference hardware, such as CPUs. Finally, some of the best-performing re-ranking methods, such as Cohere-Rerank [19], while observed in literature to be competitive with some of the best models [20], are currently only accessible via an online API.
@@ 3.1 Existing Libraries
As of the making of this work, no library providing a consistent API to access common reranking methods existed, as such, rerankers has no direct equivalent.

### v2

@@ 1 Introduction
Over time, there has been an increasing number of neural re-ranking methods. Most early neural re-ranking models were cross-encoders  [13], framing the task as a binary sentence classification task and using a pre-trained encoder model such as BERT [4]. However, many new methods have appeared, such as MonoT5 [12], leveraging sequence-to-sequence models such as T5c̃itet5 rather than BERT-like models, or ColBERT-based methods, which repurpose late-interaction retrieval mdoels such as ColBERTv2 [20] as re-ranking models. Other work also explores different paradigms, such as listwise re-ranking using sequence-to-sequence models, where the model is not tasked with outputting a score for a [query, document] pairs, but rather, to directly output a list of documents ordered by their relevance to the given query [23]. Work such as FlashRank [3], has focused on optimised methods to speed up re-ranking when performed on weaker inference hardware, such as CPUs. Finally, some of the best-performing re-ranking methods, such as Cohere-Rerank [21], while observed in literature to be competitive with some of the best models [22], are currently only accessible via an online API.
@@ 3.1 Existing Libraries
As of the making of this work, no library providing a consistent API to access common reranking methods existed, as such, rerankers has no direct equivalent. Extensive information retrieval frameworks such as Terrier [14] and Anserini [9] do integrate re-ranking capabilities, however, these aim to be much more full-featured retrieval platforms, and take a fundamentally different development approach to rerankers’ low footprint approach.